OpenStudy (dls):

The acceleration "a" in m/s^2 of a particle is given by 3t^2+2t+2,where "t" is the time.If the partile starts out with velocity V=2m/s at t=0,then,the velocity at the end of 2 second is?

OpenStudy (vincent-lyon.fr):

Do you know how to transform expression of acceleration to velocity?

OpenStudy (dls):

no..

OpenStudy (vincent-lyon.fr):

Do you have definitions of displacement, velocity, and acceleration?

OpenStudy (dls):

yes

OpenStudy (vincent-lyon.fr):

So, what is the definition of acceleration?

OpenStudy (dls):

displacement=shortest distance between 2 initial points velocity=rate of change of speed per unit time acceleration=rate of change of velocity per unit time

OpenStudy (vincent-lyon.fr):

"acceleration=rate of change of velocity per unit time" so, mathematically, acceleration is obtained by derivation of velocity and velocity by the opposite transformation, i.e. integration.

OpenStudy (dls):

i guessed that correct in my mind !

OpenStudy (vincent-lyon.fr):

So, can you integrate \(3t^2+2t+2\) ?

OpenStudy (dls):

t^3 + t^2 + 2t

OpenStudy (dls):

8+4+4 16m/s is the velocity

OpenStudy (vincent-lyon.fr):

Almost! You have just forgotten initial velocity is not zero.

OpenStudy (dls):

:o

OpenStudy (dls):

why?

OpenStudy (vincent-lyon.fr):

Read you qn again.

OpenStudy (dls):

oh sorry its 2!

OpenStudy (dls):

v=2+18t 16=2+18t 14=18t?

OpenStudy (vincent-lyon.fr):

????? I said what you did was OK, except you had forgotten that at t=0 , v=2

OpenStudy (dls):

wait what :/ im confused,where did i go wrong!

OpenStudy (vincent-lyon.fr):

You wrote: \(\large v(t) = t^3 + t^2 + 2t\) you have just forgotten initial velocity (constant of integration): You should have written: \(\large v(t) = t^3 + t^2 + 2t+V\)

OpenStudy (dls):

ah! lol sorry :/ so..velocity=18m/s

OpenStudy (vincent-lyon.fr):

yep!

OpenStudy (dls):

18=2+14t 16=14t lol still getting the same

OpenStudy (vincent-lyon.fr):

What are these equations? Haven't you realised your problem is now solved!

OpenStudy (dls):

thanks!